\begin{thebibliography}{99}

\bibitem{DYZ}Y. Duan, H. Yu, C. Zhang, Quantitative unique continuation and observability on an equidistributed set for the diffusion equation in $\mathbb{R}^N$, arXiv preprint, arXiv: 2108.04540, (2021).

\bibitem{NHM} H. Nguyen, A dependence of the cost of fast controls for the heat equation on the support of initial datum. SIAM J. Control Optim. 60 (2022), 530-544.
\bibitem{Brezis} H. Brezis, Functional Analysis, Sobolev Space and Partial Differential Equations, Springer, New York, 2011.
\bibitem{Apraiz-Escauriaza22} J. Apraiz, L. Escauriaza, Null-control and measurable sets, ESAIM: COCV 19 (2013), 239-254.

\bibitem{Apraiz-Escauriaza-Wang-Zhang} J. Apraiz, L. Escauriaza, G. Wang, C. Zhang, Observability inequalities and measurable sets, J. Eur. Math. Soc.  16 (2014), 2433-2475.
\bibitem{Escauriaza-Montaner-Zhang} L. Escauriaza, S. Montaner, C. Zhang,
	Observation from measurable sets for parabolic analytic evolutions and applications, J. Math. Pures Appl. 9 (5) (2015), 837-867.

\bibitem{Escauriaza-Montaner-Zhang2} L. Escauriaza, S. Montaner, C. Zhang, Analyticity of solutions to parabolic evolutions and
applications, SIAM J. Math. Anal. 49 (2017), 4064-4092.


\bibitem{FI} A. V. Fursikov, O. Y. Imanuvilov, Controllability of Evolution Equations, Seoul National University, 1996.

\bibitem{LEBEAU2} G. Lebeau, D. Jerson, Nodal sets of sums of eigenfunctions. Harmonic analysis and partial differential equations. Chicago Lectures
in Mathematics, University Chicago Press, Chicago, IL, 1999.

\bibitem{LEBEAU1} G. Lebeau, L. Robbiano,  Contr$\hat{\mathrm{o}}$le exact de l'$\acute{\mathrm{e}}$quation de la
chaleur. Communications in Partial Differential Equations, 20(1995), 335-356.

\bibitem{LEBEAU3} G. Lebeau, E. Zuazua, Null controllability of a system of linear
thermoelasticity. Archive for Rational Mechanics and Analysis, 141(1998), 297-329.

\bibitem{Phung-Wang-2013} K. D. Phung, G. Wang, An observability estimate for parabolic equations from
a measurable set in time and its applications, J. Eur. Math. Soc.  15 (2013), 681-703.

\bibitem{Phung-Wang-Zhang} K. D. Phung, L. Wang, C. Zhang, Bang-bang property for time optimal control
of semilinear heat equation, Ann. Inst. H. Poincar\'e Anal. Non Lin\'eaire 31 (2014), 477-499.
\bibitem{wang-zhang1} G. Wang, C. Zhang, Observability inequalities from measurable sets for some evolution equations, SIAM J. Control Optim. 55 (2017), 1862-1886.

\bibitem{MZ} S. Micu, E. Zuazua, On the lack of null-controllability of the heat equation on the half-line, Trans. Amer. Math. Soc., 353 (2001), 1635-1659.

\bibitem{MZb} S. Micu, E. Zuazua, On the lack of null-controllability of the heat equation on the half space, Port. Math. 58 (2001), 1-24.

\bibitem{M05a} L. Miller, On the null-controllability of the heat equation in unbounded domains, Bull. Sci.  math. 129 (2005), 175-185.

\bibitem{CMZ} V. R. Cabanillas, S. B. de Menezes, E. Zuazua, Null controllability in unbounded domains for the semilinear heat equation with nonlinearities involving gradient terms, J. Optim. Theory  App. 110 (2001), 245-264.
\bibitem{B} V. Barbu, Exact null internal controllability for the heat equation on unbounded convex domains, ESAIM: COCV 20 (2014), 222-235.

\bibitem{CMV} P. Cannarsa, P.  Martinez, J.  Vancostenoble, Null controllability of the heat equation in unbounded domains by a finite measure control region, ESAIM: COCV 10 (2004), 381-408.
\bibitem{Gde} M. Gonz\'alez-Burgos, L. de Teresa, Some results on controllability for linear and nonlinear heat equations in unbounded domains, Adv. Diff. Equ., 12 (2007), 1201-1240.

\bibitem{RM} J. Le Rousseau, I. Moyano, Null-controllability of the Kolmogorov equation in the whole phase space, J. Differ. Equ. 260 (2016), 3193-3233.
\bibitem{Z16} Y. Zhang,  Unique continuation estimates for the Kolmogorov equation in the whole space,
C. R. Math. Acad. Sci. Paris., 354 (2016), 389-393.
\bibitem{WangZhangZhang} G. Wang, M. Wang, C. Zhang, Y. Zhang,  Observable set, observability, interpolation inequality and spectral inequality for the heat equation in $\mathbb{R}^N$,  J. Math. Pures  Appl. 126 (2019), 144-194.

\bibitem{EV17} M. Egidi, I. Veseli\'{c}, Sharp geometric condition for null-controllability of the heat equation on
$\mathbb R^d$ and consistent estimates on the control cost, Arch. Math. (Basel) 111 (2018), 85-99.

\bibitem{nttv} I. Naki$\acute{\mathrm{c}}$, M. T$\ddot{\mathrm{a}}$ufer,  M. Tautenhahn,  I. Veseli$\acute{\mathrm{c}}$, Sharp estimates and homogenization of the control cost of the heat equation on large domains, ESAIM: COCV 26 (2020) 26-54.
\bibitem{Logunov-Malinnikova} A. Logunov, E. Malinnikova, Quantitative propagation of smallness for solutions of elliptic
equations, Proceedings of the International Congress of Mathematicians-Rio de Janeiro 2018. Vol. III. Invited lectures, 2391-2411, World Sci. Publ., Hackensack, NJ, 2018.









\bibitem{EV19} M. Egidi, A. Seelmann, The reflection principle in the control problem of the heat equation, J. Dyn. Control Syst. 28 (2022), 635-655.




%\bibitem{WangGengsheng} G. Wang, $L^{\infty}$-null controllability for the heat equation
%and its consequences for the time optimal control problem. SIAM J. Control Optim.,
%47 (2008), 1701-1720.



\end{thebibliography}

\documentclass[smallextended,referee,envcountsect]{svjour3}       
\usepackage{amsmath}
\smartqed  
\usepackage{graphicx}
\usepackage{mathptmx}      
\usepackage{latexsym}
\usepackage{bm}
\usepackage{amssymb}
\usepackage{amsfonts}

\begin{document}

\title{A Fast Observability for Diffusion Equations in $\mathbb R^N$
}


\author{Yueliang Duan \and Can Zhang }


\institute{Yueliang Duan \at
Department of Mathematics, Shantou University\\
Shantou 515063, China\\
ylduan@stu.edu.cn
\and
Can Zhang, Corresponding author \at
School of Mathematics and Statistics, Wuhan University \\
Wuhan 430072, China\\
canzhang@whu.edu.cn\\
}

\date{Received: date / Accepted: date}

\maketitle

\begin{abstract}
Given an equidistributed set in the whole Euclidean space, we have established  in \cite{DYZ} that there exists a constant positive  $C$ such that the observability inequality of diffusion equations holds for all $T\in]0,1[$, with an observability cost  being of the form $Ce^{C/T}$.
In this paper, for any small constant $\varepsilon>0$, we prove that
there exists  a nontrivial equidistributed set (in the sense that whose complementary set is unbounded), so that the above observability cost can be
improved to a fast form of $Ce^{\varepsilon/T}$ for certain constant $C>0$.
The proof is based on the strategy used in  \cite{DYZ}, as well as
an interpolation inequality for gradients of solutions to elliptic equations
obtained recently in \cite{NHM}.
\end{abstract}
\keywords{Observability \and Unbounded domain \and Equidistributed set \and Unique continuation}
\subclass{35B60  \and 35K10 \and 93B07  \and 93C20 }

\section{Introduction and main result}
\label{sec_intro}
Let $N$ be a positive integer, and let $\mathbb{R}^+=]0,+\infty[$.  We consider the diffusion equation in the whole Euclidean space
\begin{equation}\label{yu-6-24-1}
\left\{
\begin{array}{lll}
\partial_t u(x,t)-\sum\limits_{j,k=1}^N\partial_j\left(a_{jk}(x)\partial_ku(x,t)\right)=0&\mbox{in}\;\;\mathbb R^N\times\mathbb R^+,\\
		u(x,0)=u_0(x) &\mbox{in}\;\;\mathbb R^N,
\end{array}\right.
\end{equation}
where $u_0\in L^2(\mathbb R^N)$, $a_{jk}(x)=a_{kj}(x)$ for all $j,k=1,\cdots,N$, and for all  $x\in\mathbb R^N$.  Assume that the diffusion coefficient matrix $A(\cdot)=(a_{jk}(\cdot))_{j,k=1}^N$ is globally Lipschitz continuous,  and that it satisfies the uniform ellipticity condition.
That is to say, there is a constant $\lambda\geq1$ such that
\begin{equation*}\label{yu-11-28-2}
	|a_{jk}(x)-a_{jk}(y)|\leq \lambda|x-y|\quad\text{and}\quad
    \lambda^{-1}|\xi|^2\leq \sum_{j,k=1}^N a_{jk}(x)\xi_j\xi_k \leq \lambda|\xi|^2,
\end{equation*}
for all $j,k=1,\cdots,N$, $x, y\in\mathbb{R}^N$, and for all $\xi=(\xi_1,\xi_2,\cdots,\xi_N)\in\mathbb{R}^N$.
According to \cite[Theorem 10.9]{Brezis}, Equation \eqref{yu-6-24-1} has a unique solution
$u\in L^2(\mathbb{R}^+; H^1(\mathbb{R}^N))\cap C([0,+\infty[; L^2(\mathbb{R}^N))$. Moreover, if the initial datum $u_0\in H^1(\mathbb R^N)$, the corresponding solution
$u\in C([0,+\infty[; H^1(\mathbb{R}^N))$
and $\partial_t u\in L^2(\mathbb{R}^+; L^2(\mathbb{R}^N))$.

The observability inequality for solutions of linear parabolic equations on bounded domains has been studied extensively; see, e.g. \cite{Apraiz-Escauriaza22,Apraiz-Escauriaza-Wang-Zhang,Escauriaza-Montaner-Zhang,Escauriaza-Montaner-Zhang2,FI,LEBEAU2,LEBEAU1,LEBEAU3,Phung-Wang-2013,Phung-Wang-Zhang,wang-zhang1} and the references therein. The arguments to prove
those inequalities are usually based on the Carleman inequality method \cite{FI} and the spectral inequality method \cite{LEBEAU1}. However, the studies on the observability inequality for parabolic equations on unbounded domains are rather few.  We note that the observability inequality maybe not true in the case that the system is evolving in the whole space, and the observation set is bounded (see, e.g., \cite{MZ,MZb,M05a}). On the other hand, the authors of \cite{CMZ} showed that, for some parabolic equations in an unbounded domain $\Omega \subset \mathbb{R}^N$, the observability inequality holds true when observations are made over an open subset $\omega \subset\Omega$, with $\Omega \backslash \omega$ being bounded. The proof there is mainly based on Carleman estimates adapted from the case of bounded domains.  We refer the reader to \cite{B,CMV,Gde,RM,Z16} for similar results.

Recently, \cite{WangZhangZhang} and \cite{EV17} independently obtained that the observability inequality for the pure heat equation holds true in the whole space, if and only if  the observation is a thick set. This could be extended to the time-independent parabolic equation associated to the Schr\"odinger operator with analytic coefficients (see \cite{nttv}).
The methods utilized in these papers are all based on the spectral inequality (or uncertainty principle). But, they are not valid any more for the case that the coefficients in parabolic equations are non-analytic. To deal with the non-analytic case, by utilizing the propagation of smallness estimate for gradients of solutions to elliptic equations \cite{Logunov-Malinnikova} and the geometry of the equidistributed observation, we establish the observability inequality for the diffusion equation in the whole space in \cite{DYZ}.  Surprisingly,  by tracking the proof there,  even under the condition that the density of thick sets approaches to one, it seems that one cannot conclude that the observability constant appearing in the exponent tends to zero.

In order to state our main result, we introduce several useful notations below. We shall write $\mathbb N^+=\{1,2,3,\cdots\}$. Let
$B_R(x_0)$ stand for an open ball in $\mathbb R^N$ with the center $x_0$ and of radius $R>0$, while
let $B_R(x_0,0)$ stand for an open ball in $\mathbb R^{N+1}$ with the center $(x_0,0)\in\mathbb R^{N}\times\mathbb R$ and of radius $R>0$.
Define the cube by $Q_R(x_0):=\{v=(v_1,v_2,\cdots,v_N):|v_i-x_{0,i}|\leq R,\;\forall  i=1,2,\cdots,N\}$,  with $x_0=(x_{0,1},x_{0,2},\cdots,x_{0,N})\in\mathbb R^N$. Denote by $\mathrm{int}(Q_{r}(x))$ the interior of $Q_{r}(x)$, and by $\partial B_R(x_0)$ the boundary of $B_R(x_0)$.
We denote the usual inner product and norm in $L^2(B_R(x_0))$ by $\langle\cdot,\cdot\rangle_{L^2(B_R(x_0))}$ and $\|\cdot\|_{L^2(B_R(x_0))}$, respectively.
Write $|\omega|$ for the Lebesgue measure of an open subset $\omega\subset\mathbb R^N$.

The first aim of this paper is to establish a new quantitative estimate of unique continuation
for solutions of \eqref{yu-6-24-1}. For the sake of simplicity,
we would like to decompose $\mathbb R^N$ as a union of cubes with the same size in the following way: Let $r>0, \sigma\in]0,1[$,  and let  $\{x_i\}_{i\in\mathbb{N}^+}\subset\mathbb R^N$ to be such that
\begin{equation*}
   \mathbb{R}^{N}=\bigcup_{i\in\mathbb{N}^+}Q_{r+\sigma}(x_{i}),
\quad \text{with}\quad \mathrm{int}(Q_{r+\sigma}(x_{i}))\bigcap \mathrm{int}(Q_{r+\sigma}(x_{j}))=\emptyset
\end{equation*}
for different integers $i$ and $j$.
We call a proper subset $\omega^\sigma\subset\mathbb R^N$ as an observable set with
 the  parameter $\sigma$, if it satisfies
\begin{equation*}
   \omega^\sigma=\bigcup_{i\in\mathbb{N}^+} \omega^\sigma_i
\quad \text{with}\quad B_{r}(x_{i})\subset\omega^\sigma_{i} \subset B_{r+\sigma}(x_{i}) \quad\text{for each}\quad i\in\mathbb N^+.
\end{equation*}

We are now in position to state a H\"older-type interpolation inequality of quantitative unique continuation
for solutions of \eqref{yu-6-24-1}.
\begin{theorem}\label{yu-theorem-7-10-6}
Let $\varepsilon\in ]0,1[$. There exist constants $C=C(\varepsilon, N,\lambda)>0$, $\delta=\delta(\varepsilon,N,\lambda)$ $\in]0,1[$, and an observable region
$\omega^\sigma$, with  $\sigma=\sigma(\varepsilon, N, \lambda)>0$, such that
for any $T\in]0,1[$ and for any $u_{0}\in L^{2}(\mathbb R^N)$,
the corresponding solution $u$ of \eqref{yu-6-24-1} satisfies
\begin{equation*}\label{yu-7-10-2}
	\int_{\mathbb R^N}|u(x,T)|^2dx\leq Ce^{\frac{\varepsilon}{T}}\left(\int_{\omega^\sigma}|u(x,T)|^2dx\right)^{\delta}\left(\int_{\mathbb R^N}|u_{0}(x)|^2dx\right)^{1-\delta}.
	\end{equation*}
\end{theorem}

As a consequence of Theorem \ref{yu-theorem-7-10-6},  by the telescoping series method, we can obtain the final-state observability inequality in any short time for solutions to the equation \eqref{yu-6-24-1}.

\begin{theorem}\label{jiudu4}
Let  $\varepsilon\in ]0,1[$.
There exist a constant $C=C(\varepsilon,N,\lambda)>0$ and an observable region
$\omega^\sigma$, with a positive parameter $\sigma=\sigma(\varepsilon, N, \lambda)$,  such that for any $u_{0}\in L^{2}(\mathbb R^N)$ and for any $T\in]0,1[$, the corresponding solution $u$ of \eqref{yu-6-24-1} satisfies
	\begin{equation*}
	\int_{\mathbb R^N}|u(x,T)|^2dx\leq
	Ce^{\frac{\varepsilon}{T}}\int_{0}^T\int_{\omega^\sigma}|u(x,t)|^2dxdt.
\end{equation*}
\end{theorem}

\begin{remark}
We do not know how to capture the quantitative dependence on $\varepsilon$ for constants $C$ and $\sigma$ appearing in the above theorems. This remains an open problem for us.
\end{remark}
By the classical duality between the null controllability and observability inequality for linear parabolic equations, we have the following corollary:
\begin{corollary}
Let  $\varepsilon\in ]0,1[$.
There exist a constant $C=C(\varepsilon,N,\lambda)>0$ and a control region
$\omega^\sigma$, with a positive parameter $\sigma=\sigma(\varepsilon, N, \lambda)$,  such that
for any $T>0$ and for each $y_{0}\in L^2(\mathbb{R}^{N})$,
there is a control $f\in L^{2}(\mathbb{R}^{N}\times]0,T[)$ supported only on $\omega^\sigma\times]0,T[$, with the control cost
\begin{equation*}
\|f\|_{L^2(\mathbb R^N\times]0,T[)}\leq Ce^{\frac{\varepsilon}{T}}\|y_0\|_{L^2(\mathbb R^N)},
\end{equation*}
so that the corresponding solution to the controlled equation
\begin{equation*}\label{nullcontrol22}
\left\{ \begin{array}{lll}
\partial_t y(x,t)-\sum\limits_{j,k=1}^N\partial_j\left(a_{jk}(x)\partial_ky(x,t)\right)=f&\mbox{in}\;\;\mathbb R^N\times\mathbb ]0,T[,\\
		y(x,0)=y_0(x) &\mbox{in}\;\;\mathbb R^N,
\end{array}\right.
\end{equation*}
satisfies $y(x,T)=0$ for a.e. $x\in\mathbb R^N$.
\end{corollary}

In addition to the tools developed in \cite{DYZ},
the proof of Theorem \ref{yu-theorem-7-10-6} relies heavily on a new propagation estimate of smallness  for gradients of solutions to elliptic equations established in \cite{NHM}.  Precisely, we first prove a locally quantitative estimate of unique continuation for solutions to the diffusion equation.
Combing with the above local result and the geometry of the equidistributed observation, we then obtain a globally quantitative estimate at one time point for solutions of the diffusion equation. We finally apply the telescoping series method to prove Theorem \ref{jiudu4}.

The rest of this paper is organized as follows. In Section \ref{sc2}, we quote three auxiliary lemmas, which are basically taken from \cite{DYZ,NHM}.  Sections \ref{kaodu3} and \ref{finalproof} prove Theorems \ref{yu-theorem-7-10-6} and \ref{jiudu4}, respectively.
\section{Auxiliary lemmas}\label{sc2}
\subsection{Exponential decay}\label{kaodu1}
As this paper uses the same strategies as in \cite{DYZ}, we make extensive use of some of their results.
For the reader convenience, we will state some of the most relevant ones here. Statements made in this subsection will not be proven.
Let $0<\tau<T$ be arbitrarily fixed, and let $\xi\in C^\infty(\mathbb{R}^+;[0,1])$ be a cutoff function satisfying
\begin{equation}\label{yu-6-6-6}
\begin{cases}
	\xi= 1 &\mbox{in}\;\;[0,\tau],\\
	\xi=0 &\mbox{in}\;\; [T,+\infty[,\\
	|\xi_t|\leq \frac{C}{T-\tau}&\mbox{in}\;\;]\tau,T[,
\end{cases}
\end{equation}
where the positive  constant $C$ is independent  of $\tau$ and $T$. For each $R>0$ and $x_0\in\mathbb R^N$, we assume that  $v\in C^2(B_{R}(x_0)\times\mathbb{R}^+)$ solves the initial-boundary problem
    \begin{equation}\label{yu-11-29-4}
\begin{cases}
    \partial_t v(x,t)-\sum\limits_{j,k=1}^N\partial_j\left(a_{jk}(x)\partial_kv(x,t)\right)=0&\mbox{in}\;\;B_{R}(x_0)\times\mathbb{R}^+,\\
    v=\xi u&\mbox{on}\;\;\partial B_R(x_0)\times\mathbb{R}^+,\\
    v(\cdot,0)=0&\mbox{in}\;\; B_R(x_0),
\end{cases}
\end{equation}
where $u$ satisfies  \eqref{yu-6-24-1} with the initial value $u_0\in H^1(\mathbb{R}^N)$ and $\xi$ verifies \eqref{yu-6-6-6}.

By \cite[Lemma 2.1]{DYZ}, we have the following  exponential decay result.
\begin{lemma}\label{yu-lemma-6-10-1}
There exist positive constants $C_{1}=C_{1}(\lambda)$ and $C_{2}=C_{2}(\lambda,R,N)$ such that
\begin{equation}\label{yu-6-18-1}
 	\|v(\cdot,t)\|_{H^1(B_R(x_0))}\leq C_{1}\left(1+\frac{1}{T}\right)^{\frac{1}{2}}
 e^{\frac{C_{1}T}{T-\tau}-C_{2}(t-T)^+}F(R)\quad\text{for all}\;\; t\in\mathbb{R}^+,
 \end{equation}
where $(t-T)^+:=\max\{0,t-T\}$ and
	$F(R):=\sup\limits_{s\in[0,T]}\|u(\cdot,s)\|_{H^1(B_R(x_0))}$.
\end{lemma}

Define
\begin{equation*}\label{yu-6-18-5}
	\tilde{v}(\cdot,t)=
\begin{cases}
	v(\cdot,t)&\mbox{if}\;\;t\geq 0,\\
	0&\mbox{if}\;\;t<0,
\end{cases}
\end{equation*}
where $v$ is the solution of \eqref{yu-11-29-4}.
By Lemma \ref{yu-lemma-6-10-1}, we see the that  Fourier transform
of $\tilde{v}(x,\cdot)$ with respect to the time variable is meaningful, i.e.,
\begin{equation*}\label{yu-6-18-6}
	\hat{v}(x,\mu)=\int_{\mathbb{R}}e^{-i\mu t}\tilde{v}(x,t)dt\quad\text{for}\;\;(x,\mu)\in B_R(x_0)\times\mathbb R
\end{equation*}
is well defined.  By \cite[Lemma 2.2]{DYZ}, we moreover have
\begin{lemma}\label{yu-lemma-6-18-1}
There exist positive constants $C_{3}=C_{3}(\lambda,N)$ and $C_{4}=C_{4}(\lambda,N)$ such that  for each  $\mu\in\mathbb{R}$,
the following two estimates hold:
\begin{equation}\label{yu-6-23-5}
	\|\nabla \hat{v}(\cdot,\mu)\|^{2}_{L^2(B_r(x_0))}\leq \frac{C_{3}}{(R-2r)^{2}}\|\hat{v}(\cdot,\mu)\|^{2}_{L^2(B_{\frac{R}{2}}(x_0))}\quad \text{for all} \;\;0<r <\frac{R}{2},
\end{equation}
and
\begin{equation}\label{yu-6-22-16}
	\|\hat{v}(\cdot,\mu)\|_{L^2(B_{\frac{R}{2}}(x_0))}\leq C_{4}
	\left(T+R^{2}\right)e^{\frac{C_{1}T}{T-\tau}-\frac{\sqrt{|\mu|}R}{C_{5}}}F(R),
\end{equation}
where $C_{5}:=4e\sqrt{8\pi^2\sqrt{2+\lambda^{6}}}$, while $C_1>0$ and $F(R)$ are given in Lemma \ref{yu-lemma-6-10-1}.
\end{lemma}

\subsection{Stability estimate for elliptic equations}\label{yu-section-7-26-3}
Let $0<r<r'<R<+\infty$.  Set
\begin{equation*}\label{yuZZZ}
	D_{s}:=\left\{X=(x,x_{N+1})\in B_{r'}(x_0,0); dist(X, B_{r}(x_{0})\times\{0\})<s\right\} \;\; \mbox{for}\;\; s>0.
\end{equation*}
Let $g\in H^1(B_R(x_0,0))$ be a solution of
\begin{equation}\label{yu-6-23-9}
\begin{cases}
	\sum\limits_{j,k=1}^N\partial_j\left(a_{jk}(x)\partial_kg(x,x_{N+1})\right)+g_{x_{N+1}x_{N+1}}=0&\;\;\;\text{in}\;\; B_R(x_0,0),\\
	g(x,0)=0&\;\;\;\text{in}\;\; B_R(x_0),\\
	g_{x_{N+1}}(x,0)=z(x)&\;\;\;\text{in}\;\; B_R(x_0),
\end{cases}
\end{equation}
	where $z\in H^1(B_R(x_0))$.

In order to give the proof of our main result, we need the following auxiliary lemma, which is adapted from \cite[Lemma 2.1]{NHM}.
	
\begin{lemma}\label{auxiliary-lemma}

Given $\alpha\in]0,1[$. There exists $C_{6}=C_{6}(r,r',\alpha,\lambda)>0$ and  $\delta=\delta(r,r',$ $\alpha,\lambda)$ $\in]0,1[$  such that
\begin{equation}\label{2022-2.7}
	\|g\|_{H^1(D_{\delta})}\leq C_{6}\|g\|_{H^1(B_{r'}(x_0,0))}^{\alpha}\|z\|^{1-\alpha}_{L^2(B_{r}(x_{0}))}.
\end{equation}
\end{lemma}	
{\it Proof} Given $\alpha\in]0,1[$, by applying Lemma 2.1 in \cite{NHM} to $g$, there exist two constants $\delta=\delta(\alpha, r', r, \lambda)\in]0,1[$ and $C_{6}=C_{6}(\alpha, r', r, \lambda)>0$ such that
\begin{equation*}\label{2022-01}
	\|g\|_{H^1(D_{\delta})}\leq C_{6}\left(\|g\|_{H^{1/2}(B_{r}(x_0)\times\{0\})}+\|\partial_{x_{N+1}}g\|_{H^{-1/2}(B_{r}(x_0)\times\{0\})}\right)^{1-\alpha}\|g\|_{H^1(B_{r'}(x_0,0))}^{\alpha}.
\end{equation*}
This, together with \eqref{yu-6-23-9}, leads to \eqref{2022-2.7}.
\qed
\begin{remark}\label{auxiliary-lemma-remark}
The idea of proving Lemma \ref{auxiliary-lemma} comes from \cite[Lemma 2.1 and Proposition 1.2]{NHM}.
Indeed, we just slightly modified the presentation of Lemma 2.1 in \cite{NHM}, by changing
$\Omega\times]-1,1[$  and $\omega$ into $B_{r'}(x_0,0)$ and $B_{r}(x_0)$, respectively.
\end{remark}

\section{Proof of Theorem \ref{yu-theorem-7-10-6}}\label{kaodu3}
In order to give the proof of Theorem \ref{yu-theorem-7-10-6}, we need  a local interpolation inequality (i.e., Lemma \ref{lemma-2A2}). It is worth mentioning that the interpolation inequality  established in Lemma \ref{lemma-2A2} is similar to that obtained in \cite[Lemma 3.1]{DYZ}.
Compared  with them, however, there are two main differences: $(i)$ here $\alpha\in]0,1[$ is arbitrary, rather than there being some fixed value; $(ii)$ the observed region here is an open set, not a measurable set. Unfortunately, we currently have no good way to solve the question for the case that  the observation domain is a measurable set.

Denoting $\mathcal{A}(\cdot):=-\mbox{div}(A(x)\nabla(\cdot))$ with domain $$D(\mathcal{A}):=H_{0}^1(B_R(x_0))\cap H^2(B_R(x_0)),$$
we claim that
	there is a generic constant $C=C(\lambda,N)>0$ such that
 \begin{equation}\label{yu-6-7-9}
 	\langle\mathcal{A}f,f\rangle_{{L}^2(B_R(x_0))}\geq CR^{-2}\|f\|^2_{L^2(B_R(x_0))}
	\;\;\mbox{for each}\;\;f\in D(\mathcal{A}).
 \end{equation}
Using the Poincar\'e inequality
\begin{equation*}\label{yu-11-30-b-1}
    \int_{B_R(x_0)}|f|^2dx\leq \left(\frac{2R}{N}\right)^2\int_{B_R(x_0)}|\nabla f|^2dx\;\;\mbox{for each}\;\;f\in H_0^1(B_R(x_0)),
\end{equation*}	
we have that
\begin{equation*}\label{yu-10-12-1}
	\langle\mathcal{A}f,f\rangle_{{L}^2(B_R(x_0))}\geq\lambda^{-1}\int_{B_R(x_0)}
	|\nabla f|^2dx\geq \lambda^{-1}\left(\frac{N}{2R}\right)^{2}
	\int_{B_R(x_0)}
	|f|^2dx.
\end{equation*}

    As a consequence of \eqref{yu-6-7-9}, we see that the inverse of $(\mathcal A, D(\mathcal{A}))$ is positive, self-adjoint and compact in $L^2(B_R(x_0))$.
By the spectral theorem for compact self-adjoint operators,	there are  eigenvalues
	$\{\mu_i\}_{i\in\mathbb{N}^+}\subset \mathbb{R}^+$ and eigenfunctions
	$\{f_i\}_{i\in\mathbb{N}^+}\subset$ $ H_0^1(B_R$ $(x_0))$, which make up an orthogonal basis of $L^2(B_R(x_0))$,
such that
 \begin{equation}\label{yu-6-7-10}
 \begin{cases}
 	-\mathcal{A}f_i=\mu_if_i\;\;\mbox{and}\;\;\|f_i\|_{{L}^2(B_R(x_0))}=1&\mbox{for each}\;\;i\in\mathbb{N}^+,\\
CR^{-2}\leq \mu_1\leq \mu_2\leq \cdots
         \leq \mu_i\to+\infty&\mbox{as}\;\;i\to+\infty.
 \end{cases}
 \end{equation}

\begin{lemma}\label{lemma-2A2}
Let $T>0$, $0<r<+\infty$ and $x_0\in \mathbb{R}^{N}$. Given $\alpha\in]0,1[$. There exist constants $C_{7}=C_{7}(r,N,\lambda,\alpha)>0$ and  $\sigma=\sigma(r,N,\lambda,\alpha)\in]0,1[$ such that for each
$\tau\in]0,T/2[$, the corresponding solution  $u$ of \eqref{yu-6-24-1} with the initial value $u_{0}\in H^{1}(\mathbb R^N)$ satisfies
\begin{eqnarray*}\label{yu-7-10-2}
	&&\|u(\cdot,\tau)\|_{L^2(B_{r+\delta}(x_0))}\\
&\leq& C_{7} \left[T^{2}e^{\frac{C_{1}T}{T-\tau}}+e^{\frac{C_{7}}{\tau}}\right]^{\frac{1-\alpha}{2}}\|u(\cdot,\tau)\|_{L^2(B_{ r}(x_0))}^\alpha
	\left(\sup_{s\in[0,T]}\|u(\cdot,s)\|_{H^1(B_{\varrho r}(x_0))}\right)^{1-\alpha},
	\end{eqnarray*}
where $\varrho:=4C_{5}$, the constant $C_1>0$ is given by Lemma \ref{yu-lemma-6-10-1} and the constant $C_5$ is given by Lemma \ref{yu-lemma-6-18-1}.
\end{lemma}
{\it Proof} Let $u$ be the solution to the equation (\ref{yu-6-24-1})  with the initial value $u_{0}\in H^{1}(\mathbb R^N)$ and $R:=\varrho r$.
Let $u_1$ and $u_2$ be accordingly the solutions to
\begin{equation*}\label{yu-7-4-4}
\begin{cases}
	\partial_tu_{1}+\mathcal{A}u_{1}=0&\mbox{in} \;B_R(x_0)\times]0,2T[,\\
	u_1=u&\mbox{on}\;\;\partial B_R(x_0)\times]0,2T[,\\
	u_1(\cdot,0)=0 &\mbox{in}\;\;B_{R}(x_0)
\end{cases}
\end{equation*}	
	and
\begin{equation*}\label{yu-7-4-5}
\begin{cases}
	\partial_tu_{2}+\mathcal{A}u_{2}=0&\mbox{in}\;\;  B_R(x_0)\times]0,2T[,\\
	u_2=0&\mbox{on}\;\;\partial B_R(x_0)\times]0,2T[,\\
	u_2(\cdot,0)=u_0&\mbox{in}\;\; B_{R}(x_0).
\end{cases}
\end{equation*}
Obviously, $u=u_1+u_2$ in $B_R(x_0)\times[0,2T]$.
By the standard energy estimate for solutions of parabolic equations, we have that
\begin{equation}\label{yu-7-4-7}
	\sup_{t\in[0,T]}\|u_2(\cdot,t)\|_{H^1(B_R(x_0))}\leq C\|u_0\|_{H^1(B_R(x_0))}
\end{equation}
with $C=C(N,\lambda)>0.$

Arbitrarily fix  $\tau\in]0,T/2[$. Let $v_1$ be the solution to
 \begin{equation*}\label{yu-11-29-4-jia}
\begin{cases}
    \partial_tv_1+\mathcal{A}v_{1}=0&\mbox{in}\;\;B_{R}(x_0)\times\mathbb{R}^+,\\
    v_1=\xi u&\mbox{on}\;\;\partial B_R(x_0)\times\mathbb{R}^+,\\
    v_1(\cdot,0)=0&\mbox{in}\;\; B_R(x_0),
\end{cases}
\end{equation*}
where $\xi$ is given by \eqref{yu-6-6-6}.
It is clear that
	 $u=v_1+u_2$ in $B_R(x_0)\times[0,\tau]$.
	 We extend $v_1$ to $\mathbb R^-\times B_R(x_0)$ by zero, and still denote it by the same way.
Define
\begin{equation*}\label{yu-6-18-6jia}
	\hat{v}_1(x,\mu)=\int_{\mathbb{R}}e^{-i\mu t}v_1(x,t)dt
	\quad\text{for}\;\;(x,\mu)\in B_R(x_0)\times\mathbb R.
\end{equation*}
According to  Lemma \ref{yu-lemma-6-10-1},  $\hat{v}_1$ is well defined.

Let $\kappa:=\sqrt{2}/C_{5}$ with $C_{5}$ given in Lemma
	\ref{yu-lemma-6-18-1}. We define $V=V_1+V_2\quad\mbox{in}\;\; B_R(x_0)\times]-\kappa R,\kappa R[
$
 with
 \begin{equation}\label{yu-6-23-6jia}
	V_1(x,y)=\frac{1}{2\pi}\int_{\mathbb{R}}e^{i\tau\mu}\hat{v}_1(x,\mu)
	\frac{\sinh(\sqrt{-i\mu}y)}{\sqrt{-i\mu}}d\mu\quad\mbox{in}\;B_R(x_0)\times]-\kappa R,\kappa R[,
\end{equation}
\begin{equation}\label{yu-7-5-7}
	V_2(x,y)=\sum_{k=1}^\infty\alpha_ke^{-\mu_k\tau}f_k(x)\frac{\sinh(\sqrt{\mu_k}y)}
	{\sqrt{\mu_k}}\;\;\mbox{in}\;\;B_R(x_0)\times]-\kappa R,\kappa R[
\end{equation}
where $\{\mu_k\}_{k=1}^\infty$, $\{f_k\}_{k=1}^{\infty}$ are  given by $(\ref{yu-6-7-10})$, and
$\alpha_k=\langle u_2(\cdot,0),f_k\rangle_{L^2(B_R(x_0))}$.
Note from
Lemma \ref{yu-lemma-6-18-1} that $V_1$ is also well defined.
One can readily check that
\begin{equation}\label{yu-7-5-9}
\begin{cases}
	\mbox{div}(A(x)\nabla V(x,y))+V_{yy}(x,y)=0&\mbox{in}\;\;
	B_{\frac{R}{2}}(x_0)\times]-\kappa R,\kappa R[,\\
	V(x,0)=0&\mbox{in}\;\;B_{\frac{R}{2}}(x_0),\\
	V_y(x,0)=u(x,\tau)&\mbox{in}\;\;B_{\frac{R}{2}}(x_0).
\end{cases}
\end{equation}
 Let $0<r<r'<\frac{\kappa R}{2}<+\infty$ and $\alpha\in]0,1[$. By Lemma \ref{auxiliary-lemma}, we obtain that  there exists $C=C(r,r',\alpha,\lambda)>0$ and  $\delta=\delta(r,r',\alpha,\lambda)\in]0,1[$  such that
{\begin{equation}\label{yu-7-4-2}
	\|V\|_{H^1(D_{\delta})}\leq C\|V\|_{H^1(B_{r'}(x_0,0))}^{\alpha}\|u(\cdot,\tau)\|^{1-\alpha}_{L^2(B_{r}(x_0))}.
\end{equation}}
By the interior estimate, there is a constant $C=C(N)>0$ such that
\begin{equation}\label{yu-7-4-3}
	\|V\|_{H^1(B_{r'}(x_0,0))}\leq Cr'^{-1} \|V\|_{L^2(B_{2r'}(x_0,0))}.
\end{equation}
Hence, it follows from  (\ref{yu-7-4-2}) and (\ref{yu-7-4-3}) that
\begin{equation}\label{yu-7-4-1}
	\|V\|_{H^1(D_{\delta})}\leq C\|V\|_{L^2(B_{2r'}(x_0,0))}^{\alpha}\|u(\cdot,\tau)\|^{1-\alpha}_{L^2(B_{r}(x_0))}
\end{equation}
with $C=C(r',N,r,\lambda,\alpha)>0$.
Since $A(x)$ is Lipschitz, by the regularity theory of elliptic equations, one has
\begin{equation}\label{yu-7-4-1777}
	\|u(\cdot,\tau)\|_{L^2(B_{r+\delta}(x_0))}\leq C(r,N,\lambda)\|V\|_{H^1(D_{\delta})}.
\end{equation}
It follows from (\ref{yu-7-4-1777}) and \eqref{yu-7-4-1} that
\begin{equation}\label{yu-7-5-1}
	\|u(\cdot,\tau)\|_{L^2(B_{r+\delta}(x_0))}\leq C\|u(\cdot,\tau)\|_{L^2(B_{r}(x_0))}^{\alpha}
	\|V\|^{1-\alpha}_{L^2(B_{2r'}(x_0,0))}
\end{equation}
with $C=C(r',N,r,\lambda,\alpha)>0$.

To finish the proof, it suffices to bound the term $\|V\|_{L^2(B_{2r'}(x_0,0))}$. We will treat
$V_1$ and $V_2$ separately. On the one hand, we derive from (\ref{yu-6-23-6jia})  that
	for each $x\in B_{2r'}(x_0)\subset B_{R}(x_0)$ and $|y|<\kappa R/(4\sqrt{2})$,
\begin{eqnarray*}\label{yu-7-5-2}
	|V_1(x,y)|
	&\leq&\frac{1}{2\pi}\int_{\mathbb{R}}|\hat{v}_1(x,\mu)|\int_{-y}^y|e^{\sqrt{-i\mu}s}|dsd\mu
	\nonumber\\
	&\leq&\frac{\kappa R}{4\sqrt{2}\pi}\int_{\mathbb{R}}|\hat{v}_1(x,\mu)|e^{\frac{1}{4\sqrt{2}}\kappa\sqrt{|\mu|}R}d\mu\nonumber\\
	&\leq &\frac{\kappa R}{4\sqrt{2}\pi}\left(\int_{\mathbb{R}}|\hat{v}_1(x,\mu)|^2e^{\frac{1}{\sqrt{2}}\kappa\sqrt{|\mu|}R}d\mu\right)^{\frac{1}{2}}
	\left(\int_{\mathbb{R}}e^{-\frac{1}{2\sqrt{2}}\kappa \sqrt{|\mu|}R}d\mu\right)^{\frac{1}{2}}\nonumber\\
	&=&\frac{1}{2\pi}\left(\int_{\mathbb{R}}|\hat{v}_1(x,\mu)|^2e^{\frac{1}{\sqrt{2}}\kappa\sqrt{|\mu|}R}d\mu\right)^{\frac{1}{2}}.
\end{eqnarray*}
Hence, by Lemma \ref{yu-lemma-6-18-1}, we have for each $r'<\kappa R/(4\sqrt{2})$,
\begin{eqnarray}\label{yu-7-5-3}
	\int_{B_{2r'}(x_0,0)}|V_1|^2dxdy
	&\leq& \frac{1}{2\pi^2}C_{4}^2Re^{\frac{2C_{1}T}{T-\tau}}
	\left[T+R^{2}\right]^{2}F^2(R)
	\int_{\mathbb{R}}e^{-\frac{1}{\sqrt{2}}\kappa \sqrt{|\mu|}R}d\mu\nonumber\\
	&\leq&\frac{32e}{\pi^2}C_{5}^2C_{4}^2R^{-1}e^{\frac{2C_{1}T}{T-\tau}}
	\left[T+R^{2}\right]^{2}F^2(R).
\end{eqnarray}
On the other hand, by (\ref{yu-7-5-7}) and (\ref{yu-7-4-7}) we obtain
\begin{eqnarray}\label{yu-7-5-10}
	\int_{B_{2r'}(x_0,0)}|V_2|^2dxdy&\leq&\int_{-2r'}^{2r'}\int_{B_{R}(x_0)}|V_2|^2dxdy\\
	&\leq& \int_{-2r'}^{2r'}\sum_{k=1}^\infty\alpha_k^2e^{-2\mu_k\tau}\left|\frac{\sinh(\sqrt{\mu_k}y)}{\sqrt{\mu_k}}\right|^2dy\nonumber\\
	&\leq&\frac{e^{\frac{2r'^{2}}{\tau}}}{4C_{2}^{3/2}}R^{3}\sum_{k=1}^\infty\alpha_k^2=
	\frac{e^{\frac{2r'^{2}}{\tau}}}{4C_{2}^{3/2}}R^{3}\int_{B_R(x_0)}|u(x,0)|^2dx\nonumber\\
	&\leq& CR^{3}e^{\frac{C}{\tau}}F^2(R)
\end{eqnarray}
with $C=C(N,r',\lambda)>0$.
	Therefore, by (\ref{yu-7-5-3}) and (\ref{yu-7-5-10}), we conclude that
\begin{equation*}\label{yu-7-5-11}
	\|V\|_{L^2(B_{2r}(x_0,0))}\leq C\left[e^{\frac{C_{1}T}{T-\tau}}R^{-1}
	\left(T+R^{2}\right)^{2}+e^{\frac{C}{\tau}}R^{3}\right]^{\frac{1}{2}}F(R)
\end{equation*}
with $C=C(N,r',\lambda)>0$.
This, together with (\ref{yu-7-5-1}), means that
\begin{eqnarray*}\label{yu-7-5-12}
	\|u(\cdot,\tau)\|_{L^2(B_{r+\delta}(x_0))}
	&\leq& C\|u(\cdot,\tau)\|_{L^2(B_{r}(x_0))}^{\alpha}
	\|V\|^{1-\alpha}_{L^2(B_{2r'}(x_0,0))}\nonumber\\
&\leq& C\left[e^{\frac{C_{1}T}{T-\tau}}R^{-1}
	\left(T+R^{2}\right)^{2}+e^{\frac{C}{\tau}}R^{3}\right]^{\frac{1-\alpha}{2}}
\|u(\cdot,\tau)\|_{L^2(B_{r}(x_{0}))}^{\alpha}F(R)^{1-\alpha}
\end{eqnarray*}
with $C=C(N,r',r,\alpha,\lambda)>0$. The proof is immediately achieved by the arbitrariness of $u_0$.
\qed

\begin{lemma}\label{yu-LEMMA-7-10-6}
Let $T>0$ and $\varepsilon\in ]0,1[$. There exists $C_{8}=C_{8}(\varepsilon, N,\lambda)>0$ and fast control domain $\omega:=\omega(\varepsilon, N, \lambda)$
such that for any $u_{0}\in H^{1}(\mathbb R^N)$, the corresponding solution $u$ of \eqref{yu-6-24-1} satisfies
\begin{equation*}\label{yu-7-10-2}
	\|u(\cdot,T)\|_{L^2(\mathbb R^N)}\leq C_{8}
	\left(T^{3}+e^{\frac{C_{8}}{T}}\right)^{\frac{\varepsilon}{2}}\|u(\cdot,T)\|_{L^2(\omega)}^{1-\varepsilon}
	\left(\sup_{s\in[0,2T]}\|u(\cdot,s)\|_{H^1(\mathbb R^N)}\right)^{\varepsilon}.
	\end{equation*}
\end{lemma}
We will omit its proof, because it can be proved by the same way as one of proofs of Lemma 3.6 in \cite{DYZ}.
\vskip 5pt
\noindent\textbf{Proof of Theorem \ref{yu-theorem-7-10-6}.}
By standard energy estimates of solutions to \eqref{yu-6-24-1}, we have
\begin{equation}\label{yu-7-12-2}
	\|u(\cdot,t)\|_{H^1(\mathbb R^N)}\leq \frac{Ce^{Ct}}{\sqrt{t}}\|u_0\|_{L^2(\mathbb R^N)}
	\end{equation}
with $C=C(N,\lambda)>0$, for each $t\in]0,6T]$. Moreover, if $u_0\in H^1(\mathbb R^N)$, then
 \begin{equation}\label{yu-7-12-3}
 	\|u(\cdot,t)\|_{H^1(\mathbb R^N)}\leq Ce^{Ct}\|u_0\|_{H^1(\mathbb R^N)}\;\;\mbox{for each}\;\;
	t\in[0,6T],
 \end{equation}
where $C=C(N,\lambda)>0$.

We consider the following equation
\begin{equation*}\label{yu-7-12-12}
\begin{cases}
	v_t+\mathcal{A}v=0&\mbox{in}\;\;\mathbb R^N\times(0,4T),\\
	v(\cdot,0)=u(\cdot,\frac{T}{2})&\mbox{in}\;\;\mathbb R^N.
\end{cases}
\end{equation*}
	It is clear that $v(\cdot,t)=u(\cdot,t+T/2)$ when $t\in[0,4T]$. Moreover, by (\ref{yu-7-12-2}) we have
	$u(\cdot,T/2)\in H^1(\mathbb R^N)$,
	which means that $v\in C([0,4T];H^1(\mathbb R^N))$.
	From Lemma \ref{yu-LEMMA-7-10-6} (where $T$ and $\tilde{\omega}$ are replaced by $T/2$ and $\omega$, respectively), give $\varepsilon\in ]0,1[$, we have that
\begin{equation*}\label{yu-7-13-1}
	\left\|v\left(\cdot,\frac{T}{2}\right)\right\|_{L^2(\mathbb R^N)}\leq C\left(T^{3}+e^{\frac{C}{T}}\right)^{\frac{1-\varepsilon}{2}} \left\|v\left(\cdot,\frac{T}{2}\right)\right\|^\varepsilon_{L^2(\omega)}\left(\sup_{s\in[0,T]}\|v(\cdot,s)\|_{H^1(\mathbb R^N)}\right)^{1-\varepsilon}.
\end{equation*}
	This, along with (\ref{yu-7-12-3}), gives that
\begin{equation*}\label{yu-7-13-2}
	\left\|v\left(\cdot,\frac{T}{2}\right)\right\|_{L^2(\mathbb R^N)}\leq C\left(T^{3}+e^{\frac{C}{T}}\right)^{\frac{1-\varepsilon}{2}}e^{C(1-\varepsilon)T} \left\|v\left(\cdot,\frac{T}{2}\right)\right\|^\varepsilon_{L^2(\omega)}\|v(\cdot,0)\|^{1-\varepsilon}_{H^1(\mathbb R^N)}. \end{equation*}
Thus,
\begin{equation*}\label{yu-7-13-3}
	\left\|u\left(\cdot,T\right)\right\|_{L^2(\mathbb R^N)}\leq  Ce^{C\varepsilon(T+\frac{1}{T})} \left\|u\left(\cdot,T\right)\right\|^{1-\varepsilon}_{L^2(\omega)}\left\|u\left(\cdot,\frac{T}{2}\right)\right\|^{\varepsilon}_{H^1(\mathbb R^N)}.
\end{equation*}
	This, together with (\ref{yu-7-12-2}), completes the proof.\qed
\section{Proof of Theorem \ref{jiudu4}}\label{finalproof}

Finally, we give the proof of Theorem \ref{jiudu4}.\\

\noindent\textbf{Proof of Theorem \ref{jiudu4}.}
By Theorem \ref{yu-theorem-7-10-6} (where $r, x_{i}$ and $w_{i}$ are replaced by
$r, x_{i}$ and $w_{i}$, respectively) and Young's inequality,
for any $0\leq t_{1}<t_{2}\leq T$, we see that
 \begin{equation}\label{2019-7-9}
 \|u(t_{2})\|^{2}_{L^{2}(\mathbb{R}^{N})}\leq\varepsilon
 \|u(t_{1})\|^{2}_{L^{2}(\mathbb{R}^{N})}+
 \left(\frac{Ce^{CT}}{\varepsilon}\right)^{\alpha}e^{\frac{C\alpha}{t_{2}-t_{1}}}
 \|u(t_{2})\|^{2}_{L^{2}(\omega)} \ \ \ \mathrm{for\ each}\ \varepsilon>0,
 \end{equation}
where $\alpha:= \sigma/(1-\sigma)$.
According to Proposition 2.1 in \cite{Phung-Wang-2013},
for each $\kappa>1$ so that the sequence $\{l_{m}\}_{m\in\mathbb{N}^+}$, given by
$$
l_{m+1}=\frac{T}{\kappa^{m}}\;\;\mbox{for each}\;\;m\in\mathbb{N}^+,
$$
satisfies that
 \begin{equation}\label{3.2525251}
l_{m}-l_{m+1}\leq 3|(l_{m+1},l_{m})|.
 \end{equation}
\par
Next, let $0<l_{m+2}<l_{m+1}\leq t<l_{m}<T$. It follows from (\ref{2019-7-9}) that
\begin{equation}\label{3.2525252}
\|u(t)\|^{2}_{L^{2}(\mathbb{R}^{N})}\leq \varepsilon\|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}
+ \left(\frac{Ce^{CT}}{\varepsilon}\right)^{\alpha}
e^{\frac{C\alpha}{t-l_{m+2}}}\|u(t)\|^{2}_{L^{2}(\omega)}
 \end{equation}
for each $\varepsilon>0$. By a standard energy estimate, we have that
$$
\|u(l_{m})\|_{L^{2}(\mathbb{R}^{N})}\leq C_{27}\|u(t)\|_{L^{2}(\mathbb{R}^{N})},
$$
where $C=C(\lambda,N)\geq1.$
This, along with (\ref{3.2525252}), implies that
$$
\|u(l_{m})\|^{2}_{L^{2}(\mathbb{R}^{N})}\leq C^{2}\left(\varepsilon\|u(l_{m+2})
\|^{2}_{L^{2}(\mathbb{R}^{N})}+
\left(\frac{Ce^{CT}}{\varepsilon}\right)^{\alpha}
e^{\frac{C\alpha}{t-l_{m+2}}}\|u(t)\|^{2}_{L^{2}(\omega)}\right)
$$
for each $\varepsilon>0$, which indicates that
$$\|u(l_{m})\|^{2}_{L^{2}(\mathbb{R}^{N})}
\leq \varepsilon\|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}
+\left(\frac{Ce^{CT}}{\varepsilon}\right)^{\alpha}
e^{\frac{C\alpha}{t-l_{m+2}}}\|u(t)\|^{2}_{L^{2}(\omega)}
 $$
for each $\varepsilon>0$.

Integrating the latter inequality over $(l_{m+1},l_{m})$, we get that
\begin{equation}\label{3.2525253}
\begin{array}{lll}
 \displaystyle{}|(l_{m+1},l_{m})|\|u(l_{m})\|^{2}_{L^{2}(\mathbb{R}^{N})}
 &\leq&\displaystyle{}\varepsilon |(l_{m+1},l_{m})|\|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}\\
 &&\displaystyle{}+\left(\frac{Ce^{CT}}{\varepsilon}\right)^{\alpha}
e^{\frac{C\alpha}{l_{m+1}-l_{m+2}}}
 \int_{l_{m+1}}^{l_{m}}\|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t
\end{array}
\end{equation}
for each $\varepsilon>0$.
Here and throughout the proof of Theorem~\ref{jiudu4}. Since $l_{m}-l_{m+1}=(\kappa-1)T/\kappa^{m},$ by (\ref{3.2525253}) and (\ref{3.2525251}), we obtain that
\begin{eqnarray*}
&&\|u(l_{m})\|^{2}_{L^{2}(\mathbb{R}^{N})}\\
&\leq& \varepsilon \|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}+\frac{1}{|(l_{m+1},l_{m})|}
\left(\frac{Ce^{CT}}{\varepsilon}\right)^{\alpha}
e^{\frac{C\alpha}{l_{m+1}-l_{m+2}}}
\int_{l_{m+1}}^{l_{m}}\|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t\\
&\leq&\frac{3\kappa^{m}}{T(\kappa-1)}
\left(\frac{Ce^{CT}}{\varepsilon}\right)^{\alpha}
e^{C\alpha\left(\frac{1}{T}\frac{\kappa^{m+1}}{\kappa-1}\right)}
\int_{l_{m+1}}^{l_{m}}\|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t+
\varepsilon \|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}
 \end{eqnarray*}
 for each $\varepsilon>0$.
This yields that
\begin{equation}\label{3.2525254}
\begin{array}{lll}
 \displaystyle{}\|u(l_{m})\|^{2}_{L^{2}(\mathbb{R}^{N})}&\leq& \displaystyle{}
 \frac{1}{\varepsilon^{\alpha}}\frac{3}{\kappa}
 \frac{e^{C\alpha T}}{C\alpha}
 e^{2C\alpha\left(\frac{1}{T}
 \frac{\kappa^{m+1}}{\kappa-1}\right)}\int_{l_{m+1}}^{l_{m}}
 \|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t\displaystyle{}+
 \varepsilon \|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}
\end{array}
\end{equation}
for each $\varepsilon>0$.
Denote $d:= 2C\alpha/[T\kappa(\kappa-1)]$.
It follows from (\ref{3.2525254}) that
\begin{eqnarray*}
\varepsilon^{\alpha}e^{-d\kappa^{m+2}}\|u(l_{m})\|^{2}_{L^{2}(\mathbb{R}^{N})}
-\varepsilon^{1+\alpha}e^{-d\kappa^{m+2}}\|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}
\leq\frac{3}{\kappa}
 \frac{e^{C\alpha T}}{C\alpha}\int_{l_{m+1}}^{l_{m}}\|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t
\end{eqnarray*}
for each $\varepsilon>0$.
Choosing $\varepsilon=e^{-d\kappa^{m+2}}$ in the latter inequality, we observe that
\begin{equation}\label{3.25252555}
\begin{array}{lll}
 &&\displaystyle{}e^{-(1+\alpha)d\kappa^{m+2}}\|u(l_{m})\|^{2}_{L^{2}(\mathbb{R}^{N})}
 -e^{-(2+\alpha)d\kappa^{m+2}}\|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}\\
 &\leq&\displaystyle{}\frac{3}{\kappa}
 \frac{e^{C\alpha T}}{C\alpha}\int_{l_{m+1}}^{l_{m}}\|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t.
\end{array}
\end{equation}
Take $\kappa=\sqrt{(\alpha+2)/(\alpha+1)}$ in (\ref{3.25252555}). Then we have that
\begin{eqnarray*}
e^{-(2+\alpha)d\kappa^{m}}\|u(l_{m})\|^{2}_{L^{2}(\mathbb{R}^{N})}
-e^{-(2+\alpha)d\kappa^{m+2}}\|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}
\leq \frac{3}{\kappa}\frac{e^{C\alpha T}}{C\alpha}
\int_{l_{m+1}}^{l_{m}}\|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t.
\end{eqnarray*}
Summing the above inequality from $m=1$ to infinity give the desired result. Indeed,
\begin{eqnarray*}
&&C_{27}^{-2}e^{-(2+\alpha)d\kappa^{2}}\|u(T)\|^{2}_{L^{2}(\mathbb{R}^{N})}
\leq e^{-(2+\alpha)d\kappa^{2}}\|u(l_{2})\|^{2}_{L^{2}(\mathbb{R}^{N})}\\
&\leq&\sum_{m=1}^{+\infty}\left(e^{-(2+\alpha)d\kappa^{m}}\|u(l_{m})\|_{L^{2}(\mathbb{R}^{N})}
-e^{-(2+\alpha)d\kappa^{m+2}}\|u(l_{m+2})\|^{2}_{L^{2}(\mathbb{R}^{N})}\right)\\
&\leq& \frac{3}{\kappa}\frac{e^{C\alpha T}}{C\alpha}\sum_{m=1}^{+\infty}
\int_{l_{m+1}}^{l_{m}}\|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t
\leq \frac{3}{\kappa}\frac{e^{C\alpha T}}{C\alpha}\int_{0}^{T}\|u(t)\|^{2}_{L^{2}(\omega)}\mathrm{d}t.
 \end{eqnarray*}

In summary, we finish the proof of Theorem~\ref{jiudu4}.\qed

\begin{small}
\noindent\textbf{Acknowledgements}
The first author is supported by the National Natural Science Foundation of China under grant 12201379 and STU Scientific Research Foundation for Talents under grant NTF21044.
The second author is supported by the National Natural Science Foundation of China under grant 11971363, and by the Fundamental Research Funds for the Central Universities  under grant 2042023kf0193.
\end{small}

\end{document}